# Tint Your Models Task-wise for Improved Multi-task Model Merging

### v1

@@ B.2 Baseline Details
Appendix C More Experimental Results
Figure B: Multi-task performance when merging RoBERTa models on 8 tasks. Yellow indicates training only coefficients, green indicates training only the classifier, and blue indicates training both in our method.
C.1 NLP Tasks
We conduct experiments to investigate how the training of the classifier and merging coefficients, respectively, affects the performance of NLP tasks in Model Tinting , as presented in Table 2 of the main paper. Figure B shows the results of merging RoBERTa [31] models for 8 NLP tasks using the predictions of individual models as guidance, optimized through the cross-entropy loss. As mentioned in the main paper, evaluation metrics differ across tasks: the Matthews correlation coefficient is used for CoLA [56], the average of Pearson and Spearman correlations is applied to STS-B [4], and the accuracy is used for all other tasks [47, 11, 20, 57, 44, 13]. The approach where both the classifier and coefficients are trained simultaneously corresponds to Model Tinting . To examine the role of classifier training in NLP tasks, ablation experiments are conducted by removing specific components. The results show that training only the coefficients leads to the lowest performance while training only the classifier achieves a relatively high performance. Furthermore, training both the classifier and coefficients together demonstrated a complementary effect, achieving the highest performance.
C.2 Vision Tasks
@@
Figure C: Impact of training different task-specific layers with merging coefficients. We compare average performance for merged models with different trainable task-specific layers while other layers update only their merging coefficients.
Appendix D More Empirical Analyses
D.1 On Task-specific Layers
Layer SUN Cars RES. Euro SVH. GTS. MNI. DTD Avg.
0 66.2 71.9 92.2 99.0 94.2 96.5 99.1 77.0 87.0
1 65.1 69.9 92.4 98.8 94.2 97.4 99.0 77.6 86.8
2 64.4 69.8 92.3 99.0 93.8 97.3 98.6 77.4 86.6
3 64.3 68.3 92.2 99.1 92.8 98.3 99.1 78.2 86.5
4 64.8 69.5 93.3 98.6 94.6 98.1 99.2 78.2 87.0
5 64.6 70.7 91.3 99.0 94.8 98.3 99.1 78.8 87.1
6 67.4 70.7 94.4 98.9 92.3 98.3 99.2 79.1 87.5
7 67.1 66.8 94.4 99.7 94.8 99.2 99.3 79.2 87.6
8 66.1 69.8 93.6 99.0 94.2 98.2 99.5 78.6 87.4
9 68.2 66.5 94.6 99.6 91.9 98.5 99.2 79.5 87.3
10 68.3 73.0 94.0 99.6 92.5 97.9 99.1 79.4 88.0
11 68.2 69.0 94.2 99.4 92.3 97.5 99.0 79.3 87.4
Classifier 69.9 73.8 92.2 98.6 89.5 97.5 99.0 78.8 87.4
Late 50.7 63.4 89.3 99.6 93.4 95.9 99.3 75.5 83.4
Early 42.5 54.2 85.0 93.8 93.2 96.3 98.2 76.0 79.9
All 25.5 31.2 82.7 97.0 92.4 96.2 97.5 72.7 74.4
(a) Training layers.
@@ D.1 On Task-specific Layers
(b) Training layers with merging coefficients.
Table C: Performance details of merged ViT-B/32 models when training different layers for task-specific adjustments. We observe that consistently high performance can be achieved regardless of which task-specific single layer is trained. Note that the 0-th layer in the table refers to the first transformer layer.
Our proposed method, Model Tinting , introduces the concept of incorporating a task-specific layer to balance shared and task-specific representations during model merging. This approach allows effective adaptation while maintaining efficiency to address task conflicts in multi-task learning. Building on the results in Figure 4 of the main paper, which demonstrate the potential of training inner layers while keeping the merging coefficients fixed, we extend this analysis further.
Specifically, we explore whether performance remains robust when both merging coefficients and task-specific layers are trained jointly. ViT-B/32 is employed and initialized using Task Arithmetic-merged weights across all layers again. When a particular layer is marked as task-specific, it is split into unique versions for each task, resulting in eight task-specific layers. While these layers are trained exclusively on corresponding task data, the remaining layers update only the merging coefficients using data from all tasks.
Figure C shows that training both the merging coefficients and task-specific layers achieves performance comparable to or slightly better than training merging coefficients alongside the classifier. These findings highlight the minimal impact of task-specific layer choice on overall performance, reaffirming the robustness of our approach. Table C provides detailed task-level results, while Table C shows consistent average performance across layers, with specific layers excelling on certain datasets (e.g., the 11th layer on SUN397 [61] and the 7th layer on GTSRB [48]). Conversely, Table C demonstrates that merging coefficients adaptively balance representations across tasks, preventing biases even when other layers remain frozen.
@@ D.2 On Loss Functions
Figure D: Comparison of training with various loss functions. The yellow bars represent the case where only the coefficients are trained, while the blue bars indicate the joint training of the classifier and coefficients. The dashed line corresponds to training with ground truth labels using cross-entropy loss, which is the upper bound.
@@ D.3 Convergence Analysis
Figure F: Learning curves of our method on the 8 vision tasks across several models and learning rates.
@@
Figure G: Impact of joint training on coefficient sparsity on ViT-B/32.
@@
Figure H: Impact of joint training on coefficient sparsity on ViT-B/16.
@@
Figure I: Impact of joint training on coefficient sparsity on ViT-L/14.
D.4 Sparsity Visualization
The main paper highlights that sparsity in merging coefficients plays a critical role in reducing task conflicts during multi-task model merging as shown in Figure 3. To confirm this trend across various backbones, Figure GH and I provide the layer-wise merging coefficients for ViT-B/32, ViT-B/16 and ViT-L/14. All models are optimized using the cross-entropy loss with the predictions of individual models as guidance. In these figures, (a) represents training only the coefficients, while (b) includes joint training of the classifier and coefficients. Joint training consistently increases the proportion of near-zero coefficients (from 35.3% to 56.6% for ViT-B/16 and from 23.3% to 55.0% for ViT-L/14). Correspondingly, the average accuracy also improves (from 86.8% to 91.8% for ViT-B/16 and from 91.5% to 93.8% for ViT-L/14). These results indicate that training the classifier complements sparsity in merging coefficients, effectively reducing task conflicts and enhancing performance.

### v2

@@ B.2 Baseline Details
LiNeS [64] observes that reducing the influence of shallow layers helps prevent distortion of general representations, thereby simplifying merging coefficient selection by allowing them to increase linearly with layer depth.
@@
(a) Fixed Classifier
(b) Classifier Training with Task Arithmetic
(c) Classifier Training with AdaMerging
Figure B: Cross-task evaluations across diverse encoders and heads from different tasks. Each cell $(i,j)$ displays the accuracy when features from task $i$’s visual encoder are classified by task $j$’s classifier. The upper value in a cell shows the absolute task accuracy, and the number below (in parentheses) denotes the relative performance compared to the diagonal. (a) Base (untrained) classifiers generally deteriorate when evaluated on other tasks or using different encoders. (b) Classifiers trained with Task Arithmetic features show mostly reduced performance loss (i.e., higher relative numbers). This suggests that task-specific adjustments effectively refine the task-agnostic representations task-wise. (c) Classifiers trained with AdaMerging features also improve over (a) but perform worse than (b), highlighting the importance of a shared encoder when leveraging task-specific knowledge.
Appendix C More Experimental Results
C.1 Pilot Study
Figure 2(a) shows cross-task evaluations via individual visual encoders with their respective base classifiers; Figure 2(b) shows results from combining individual visual encoders with trained classifiers from Task Arithmetic, while Figure 2(c) presents results using the AdaMerging merged encoder. We observe consistent improvements in cross-task performance when using the classifiers trained on other tasks. For instance, integrating the GTSRB [54] classifier trained on Task Arithmetic features with the SVHN [41] encoder achieves 56.3% accuracy (Figure 2(b)), surpassing the base classifier’s 31.4% (Figure 2(a)). These results suggest that training the classifier on merged features for the GTSRB task also facilitates the alignment of individual SVHN features with the GTSRB task. This demonstrates the trained layer’s enhanced ability to refine task-specific outputs from task-agnostic representations. On the other hand, classifiers trained on AdaMerging features achieve 49.0% accuracy (Figure 2(c)), surpassing the base classifier but performing worse than the classifier trained on Task Arithmetic features. This suggests that even when task-specific layers are trained, the choice of a shared encoder can lead to suboptimal performance. Therefore, it highlights the importance of the shared encoder in effectively leveraging task-specific knowledge.
@@ C.1 Pilot Study
Figure C: Multi-task performance when merging RoBERTa models on 8 tasks. Yellow indicates training only coefficients, green indicates training only the classifier, and blue indicates training both in our method.
C.2 NLP Tasks
We conduct experiments to investigate how the training of the classifier and merging coefficients, respectively, affects the performance of NLP tasks in Model Tinting , as presented in Table 3 of the main paper. Figure C shows the results of merging RoBERTa [35] models for 8 NLP tasks using the predictions of individual models as guidance, optimized through the cross-entropy loss. As mentioned in the main paper, evaluation metrics differ across tasks: the Matthews correlation coefficient is used for CoLA [65], the average of Pearson and Spearman correlations is applied to STS-B [5], and the accuracy is used for all other tasks [53, 12, 24, 66, 49, 14]. The approach where both the classifier and coefficients are trained simultaneously corresponds to Model Tinting . To examine the role of classifier training in NLP tasks, ablation experiments are conducted by removing specific components. The results show that training only the coefficients leads to the lowest performance while training only the classifier achieves a relatively high performance. Furthermore, training both the classifier and coefficients together demonstrated a complementary effect, achieving the highest performance.
C.3 Vision Tasks
@@
Table C: Performance details of merged ViT-B/32 models when training different layers with merging coefficients for task-specific adjustments. We observe that consistently high performance can be achieved regardless of which task-specific single layer is trained.
Appendix D More Empirical Analyses
D.1 Detailed Results on Task-specific Layers
Our proposed method, Model Tinting , introduces the concept of incorporating a task-specific layer to balance shared and task-specific representations during model merging. This approach allows effective adaptation while maintaining efficiency to address task conflicts in multi-task learning. ViT-B/32 is employed and initialized using Task Arithmetic-merged weights across all layers. When a particular layer is marked as task-specific, it is split into unique versions for each task, resulting in eight task-specific layers. While these layers are trained exclusively on corresponding task data, the remaining layers update only the merging coefficients using data from all tasks. Table C presents detailed task-level results. ‘Early’ refer to layers 1 to 6-th, while ‘Late’ refer to layers 7 to 12-th. Training both the merging coefficients and task-specific layers achieves performance comparable to or slightly better than training merging coefficients alongside the classifier. These findings highlight the minimal impact of task-specific layer choice on overall performance, reaffirming the robustness of our approach. The results also show consistent average performance across layers, demonstrating that merging coefficients effectively balance representations across tasks, mitigating biases even when other layers remain frozen.
@@ D.2 On Loss Functions
Figure D: Comparison of training with various loss functions. The yellow bars represent training only the coefficients, while the blue bars indicate the joint training of the classifier and coefficients. The dashed line corresponds to training with ground truth labels using cross-entropy loss, which is the upper bound.
@@
Figure F: Learning curves of our method on the 8 vision tasks across several models and learning rates. Our method demonstrates robust and fast convergence, regardless of model size and learning rate.
@@ D.3 Convergence Analysis
D.4 Sparsity Visualization
The main paper highlights that sparsity in merging coefficients plays a critical role in reducing task conflicts during multi-task model merging. To confirm this trend across various backbones, Figure GH and I provide the layer-wise merging coefficients for ViT-B/32, ViT-B/16 and ViT-L/14. All models are optimized using the cross-entropy loss with the predictions of individual models as guidance. In these figures, (a) represents training only the coefficients, while (b) includes joint training of the classifier and coefficients. Joint training consistently increases the proportion of near-zero coefficients (from 35.3% to 56.6% for ViT-B/16 and from 23.3% to 55.0% for ViT-L/14). Correspondingly, the average accuracy also improves (from 86.8% to 91.8% for ViT-B/16 and from 91.5% to 93.8% for ViT-L/14). These results indicate that training the classifier complements sparsity in merging coefficients, effectively reducing task conflicts and enhancing performance.
@@ D.4 Sparsity Visualization
Figure G: Impact of joint training on merging coefficient sparsity on ViT-B/32.
@@
Figure H: Impact of joint training on merging coefficient sparsity on ViT-B/16.
@@
Figure I: Impact of joint training on merging coefficient sparsity on ViT-L/14.